# The power-spectrum tensor in steady-state systems and its role in quantum friction

### v1

@@ III.1 The power spectrum in quantum friction
In order to see how the previous information is relevant for understanding the physics of the system, let us consider the expression for the quantum frictional force. If the system is at zero temperature, earlier work  [73, 57, 25, 78] has shown that in the NESS the frictional force is oriented along the direction of translational invariance, i.e. $\mathbf{F}_{\rm fr}=F_{\rm fr}\mathbf{n}$ (see Fig. 1), and is of the form
@@
$F_{\rm fr}=-2\int_{0}^{\infty}\mathrm{d}\omega\int\frac{\mathrm{d}q}{2\pi}\,q\;\mathrm{Tr}\left[\underline{S}^{\sf T}(-\omega^{-}_{q},v)\underline{G}_{\Im}(q,\mathbf{R}_{a},\omega)\right].$ (23)
@@
where we indicate with ‘$\mathrm{Tr}$’ the trace over the tensors’ indices, while $q$ is the component of the wave vector along the direction of motion and $\omega^{\pm}_{q}=\omega\pm qv$ is the Doppler-shifted frequency. In addition to the dipole’s power-spectrum tensor $\underline{S}$, the interaction is described in terms of the electromagnetic Green tensor $\underline{G}$: Physically, it represents the linear electromagnetic susceptibility tensor of the system without the particle. For the aforementioned arrangement of the material bodies that comprise the system, it can be in principle obtained by solving the corresponding Maxwell equations with a source term [57, 86].
The physical interpretation of Eq. (23) identifies the l.h.s. as the net total momentum per unit of time that is transferred to or from the particle during the stochastic dynamics connected with the underlying absorption and emission processes characterizing the interaction [87, 88, 25, 54]. Specifically, $\underline{G}_{\Im}=[\underline{G}-\underline{G}^{{\dagger}}]/(2\mathrm{i})$ in Eq. (23) is connected with the electromagnetic density of states [89]. Moreover, in complete analogy to $\alpha_{\Im}$, we may relate the matrix $\underline{G}_{\Im}$ to the exchange of energy (dissipation) with the electromagnetic environment that surrounds the particle and is consequently positive semidefinite for $\omega>0$ (see also the discussion around Eq. (30)).
@@
Technically, the Green tensor represents a propagator of the electromagnetic field and as such usually depends on two spatial arguments [90, 91], a source position and a detection position. However, since in all our expressions we require the back-action of the electromagnetic field emitted by the particle onto itself (via scattering by the material bodies), source and detection positions coincide and we retain in the Green tensor only one spatial argument. In addition, given the system’s translational invariance along the particle trajectory, we carry out a Fourier transform, so that the Green tensor acquires a mixed Fourier- and real-space dependence with regards to the positional arguments and, of course, features a dependence on frequency, i.e. $\underline{G}(q,\mathbf{R}_{a},\omega)$. This expression fulfills the identities
@@
which we may, respecitively, view as consequences of the system’s reciprocity and stationarity.
@@
This would lead to an expression of the same form as provided in Eq. (20). However, due to the (system-specific) reciprocity condition, Eq. (24a), we can go one step further and find that the coefficients associated with $\bm{\underline{\Lambda}}_{\perp}$ and $\underline{L}_{n}$ actually vanish (see App. B) so that we can write
@@
$\mathrm{Tr}\left[\underline{S}^{\sf T}\underline{G}_{\Im}\right]=\mathrm{Tr}\left[\underline{\Delta}\underline{D}\right]+2\chi_{n}c_{n}-2\bm{\Omega}_{\perp}\cdot\mathbf{s}_{\perp}.$ (27a)
@@
=$\stackrel{{\scriptstyle\text{average}}}{{=}}\frac{d^{2}}{3}\mathrm{Tr}\left[\underline{G}_{\Im}\right]\;\Sigma=\frac{d^{2}}{3}\mathrm{Tr}\left[\underline{D}\right]\;\Sigma~{}.$ (27b)
@@
The last line in the above equations corresponds to the result of an average over all possible directions of the vector $\mathbf{d}$ for which the power-spectrum tensor becomes proportional to $\underline{\mathds{1}}d^{2}/3$ ($d^{2}=\mathbf{d}\cdot\mathbf{d}$)  [24, 98, 82]. When comparing the two results, we notice, as expected, that the main difference lies in the contribution arising from $\bm{\Omega}_{\perp}$, which is absent in the second model. Moreover, when $\bm{\Omega}_{\perp}$ and $\mathbf{s}_{\perp}$ have the same sign, the corresponding contribution tends to reduce the frictional force. This confirms and generalizes earlier results obtained within a specific model for the dipole’s dynamics [25, 78] (see Sec. III.2). The remaining contributions in Eqs. (III.1) are similar in both models although not identical. Interestingly, the terms proportional to $c_{n}$ directly show how the system’s symmetry (or better the lack thereof) affects the frictional force. More precisely, as soon as the total system is mirror-symmetric with respect to a plane containing the particle’s trajectory – for example when the particle moves above a planar structure – the terms proportional to $c_{n}$ in Eqs. (III.1) vanish. Also, averaging over the dipole direction in the second model, washes out this contribution which does not appear in the last line of Eq.(III.1).
@@
Another interesting consequence is that the results in Eqs. (III.1) are both equivalent to the identity
$\mathrm{Tr}\left[\underline{S}^{\sf T}(-\omega^{+}_{q},-v)\underline{G}_{\Im}(q,\mathbf{R}_{a},\omega)\right]\\ =\mathrm{Tr}\left[\underline{S}(-\omega^{+}_{q},v)\underline{G}_{\Im}(q,\mathbf{R}_{a},\omega)\right]~{}.$ (28)
@@
When used in Eq. (23) in conjunction with Eq. (24a) and after changing $q\to-q$, the previous expression allows us to show that, in general, the frictional force must always reverse its sign if $v\to-v$, i.e. $F_{\rm fr}(-v)=-F_{\rm fr}(v)$. In other words, independently from the model for the dipole’s dynamics, $F_{\rm fr}(v)$ is an odd function of the velocity. This is what is physically expected for a frictional interaction and has been confirmed by calculations with specific models [99, 100, 101, 73, 88, 54]. We want to underline, however, that the mathematical properties justifying this behavior are less trivial that one could intuitively presume. Indeed, reversing the logic and imposing that $F_{\rm fr}(v)$ is odd with respect to the velocity one would obtain Eq. (28), which can erroneously suggest that $\underline{S}^{\sf T}(\omega,-v)=\underline{S}(\omega,v)$, in contradiction with Eq. (20).
III.2 A illustrative example
@@
So far, we have refrained from considering any specific model for the microscopic dynamics of the electric dipole (vector-)operator. In order to illustrate the validity and physical meaning of our results, we now proceeed to consider a specific model which allows for an exact solution of the equations of motions. Specifically, we assume that the particle moving in vacuum parallel to a translational invariant arrangement of bodies is an atom. Given that the interaction between the atomic internal degrees of freedom and the electromagnetic fluctuations is typically weak, it is reasonable to describe the dipole operator in term of linear system associated with a single frequency $\omega_{a}$. Physically, $\omega_{a}$ would correspond to the transition frequency between the highest occupied orbital to the lowest unoccupied one. Within this linear approximation, we are going to represent $\hat{\mathbf{d}}$ in terms of a 3D isotropic quantum-mechanical oscillator  [102, 73] which fulfills the following equation of motion
@@
$\langle\hat{\mathbf{E}}_{0}(q;\mathbf{R}_{a};\omega)\hat{\mathbf{E}}^{\mathsf{T}}_{0}(\tilde{q},\mathbf{R}_{a};\tilde{\omega})\rangle\\ =\frac{8\pi^{2}\hbar}{1-e^{-\beta\hbar\omega}}\underline{G}_{\Im}(q,\mathbf{R}_{a},\omega)\delta(\omega+\tilde{\omega})\delta(q+\tilde{q})\,,$ (30)
@@
With the above results, we can explicitly check the general properties of the power-spectrum tensor which we have discussed in Sec. II and at the beginning of Sec. III. To start, we notice that $\underline{S}$ is Hermitian and in general not real-valued, as expected. It does not fulfill the FDT but, instead, fulfills Eq. (13). To see this more clearly, we consider the following identity which we can derive from Eq. (33)
@@
×$\times\underline{\alpha}(\omega,v)\underline{G}_{\Im}(q,\mathbf{R}_{a},\omega_{q}^{+})\underline{\alpha}^{{\dagger}}(\omega,v)~{}.$ (37)
Upon utilizing the properties of the polarizability and of the Green tensors mentioned above, we can show that, consistent with the discussion in Sec. II, $\underline{J}$ is Hermitian and fulfills the crossing relation. Moreover, in earlier work with our co-workers, we have demonstrated within the above framework that the local thermal equilibrium approximation displays significant drawbacks regarding the quantitative evaluation of quantum friction. Specifically, the local thermal equilibrium approximation underestimates the strength of the interaction and adversely affects the thermodynamical consistency of the physical system  [24, 25, 26].
Consider now the power-spectrum tensor. Upon noting that $(1-e^{x})^{-1}=1-(1-e^{-x})^{-1}$ and upon utilizing the expressions in Eqs. (III.1) and (36) as well as the properties of the dressed polarizability, we find
@@
With the above, we can already check that the expression in Eq. (39) has a form that is consistent with Eq. (26) where $c_{n}=0$ and $\mathbf{s}_{\perp}=s_{y}\mathbf{y}$. As discussed earlier, we may view this as being the consequence of the mirror-symmetry of the total system with respect to the $y=0$ plane. Finally, inserting in the previous equations the expression for the Green tensor given above and using Eq. (34), one can show that $\underline{S}(\omega,v)$ is given by the sum of a diagonal and an anti-diagonal matrix. This is in agreement with our general considerations about the structure of the tensor based on the symmetry of the system. Moreover, we can verify that $\underline{S}(\omega,-v)\neq\underline{S}^{\sf T}(\omega,v)$, as we should expect in terms of our general result in Eq. (20). Notice that this result is obtained despite the fact that $\underline{\alpha}(\omega,-v)=\underline{\alpha}^{\sf T}(\omega,v)$ and that from Eq. (35) $\underline{\mathcal{D}}(\omega,-v)=\underline{\mathcal{D}}^{\sf T}(\omega,v)$. One can verify that
@@
where, in agreement with the relation in Eq. (20), $f(\omega,v)$ is a real function odd in $v$. The previous expression also implicitly confirms that the diagonal elements of the power-spectrum tensor are even in $v$, while the off-diagonal terms arising from $\bm{\Omega}_{\perp}=\Omega_{y}\mathbf{y}$ are odd in the same variable.
IV Conclusions
@@
In summary, we have discussed a number of properties of the power-spectral density for a quantum system in an arbitrary steady state. This includes a large class of nonequilibrium scenarios as well as the equilibrium situation as a limiting case. We have adopted a minimalistic viewpoint and, besides stationarity, our analyses do not rely on any specific information about the microscopic dynamics of the system. Therefore, our results have a broad validity, including systems with non-Markorvian as well as nonlinear dynamics. This is particularly relevant for quantum systems in which the Markov approximation may lead to inconsistent results [4, 57].
@@
More precisely, we have focused on the steady state of a multidimensional (multivariate) stochastic process, where the dynamics is described in terms of a three-dimensional vector operator. Since the steady state may correspond to nonequilibrium dynamics for which the principle of detailed balance can be violated, equilibrium relations such as the fluctuation-dissipation theorem are not necessarily fulfilled. Nonetheless, we have been able to derive a collection of properties that characterize the power spectrum and its connection to other physical quantities such as the system’s susceptibility. In particular, we have determined the properties of quantities which can characterize the system’s steady-state nonequilibrium behavior and have defined the correction to the equilibrium fluctuation-dissipation theorem. This provides us with a solid basis to investigate and assess the validity of approaches such as the local thermal equilibrium approximation.
@@
Once additional information about the system such as symmetries and details about its structure is provided, we may obtain further relations and properties among the different quantities that characterize the system. As an illustration, we have analyzed the nonconservative viscous dynamics of a particle that moves in the quantum electromagnetic vacuum parallel to a translationally invariant arrangement of material bodies. When the drag felt by the particle during its motion is balanced by an external driving force, the system reaches a nonequilibrium steady state that is characterized by a speed $v$. Firstly, without specifying the internal microscopic dynamics of the moving object, we have derived a collection of relations that characterize the power-spectrum tensor, its dependence on the velocity $v$ and its role in determining the frictional interaction at zero temperature (quantum friction). Secondly, for this system, we have tested the general results described in main part of the manuscript, in the context of a specific model which provides the microscopic internal dynamics of a moving atom. In particular, we have modeled the atom in terms of a three-dimensional harmonic oscillator and have derived and examined the corresponding power-spectrum tensor as well as the related matrices.
@@
Our results can be useful in a large variety of systems and specifically for numerous fluctuation-induced phenomena, such as the (quantum) Brownian motion, the Casimir effect and the van der Waals force, including their nonequilbrium variants as quantum friction. Since often the information about the statistical properties of the system is or can be encoded in the power spectral density, knowing its general properties can be very helpful for understanding its behavior and to test thermodynamic consistency of its descriptions [26].
V Acknowledgements

### v2

@@ 3.1 The power spectrum in quantum friction
In order to see how the information above is relevant for understanding the physics of the system, let us consider the expression for the quantum frictional force. If the system is at zero temperature, earlier work  [67, 51, 22, 73] has shown that in the NESS the frictional force is oriented along the direction of translational invariance, i.e. $\mathbf{F}_{\rm fr}=F_{\rm fr}\mathbf{n}$ (see Fig. 2), and is of the form
@@
$F_{\rm fr}=-2\int_{0}^{\infty}\hskip-5.69054pt\mathrm{d}\omega\int\frac{\mathrm{d}q}{2\pi}\,q\;\mathrm{Tr}\left[\underline{S}^{\sf T}(-\omega^{-}_{q},v)\underline{G}_{\Im}(q,\mathbf{R}_{a},\omega)\right].$ (24)
@@
where we indicate with $\mathrm{Tr}$’ the trace over the tensors’ indices, while $q$ is the component of the electromagnetic field’s wave vector along the direction of motion and $\omega^{\pm}_{q}=\omega\pm qv$ is the Doppler-shifted frequency. In addition to the dipole’s power-spectrum tensor $\underline{S}$, the interaction is described in terms of the electromagnetic Green tensor $\underline{G}$: Physically, it represents the linear electromagnetic susceptibility tensor of the system without the particle, i.e. it exclusively characterizes the (electromagnetic) environment with which the particle interacts. For the aforementioned arrangement of the material bodies that comprise the system, $\underline{G}$ can be in principle obtained by solving the corresponding Maxwell equations with a source term [51, 81].
@@
The physical interpretation of Eq. (24) identifies the l.h.s. as the net total momentum per unit of time that is transferred to or from the particle during the stochastic dynamics connected with the underlying absorption and emission processes characterizing the interaction [82, 71, 22, 72]. Specifically, $\underline{G}_{\Im}=[\underline{G}-\underline{G}^{{\dagger}}]/(2\mathrm{i})$ in Eq. (24) is connected with the electromagnetic density of states [83]. Moreover, in complete analogy to $\alpha_{\Im}$, we may relate the matrix $\underline{G}_{\Im}$ to the exchange of energy (dissipation) with the electromagnetic environment that surrounds the particle and is consequently positive semidefinite for $\omega>0$ (see also the discussion around Eq. (31)). Technically, the Green tensor represents a propagator of the electromagnetic field and as such usually depends on two spatial arguments [84, 85], a source position and a detection position. However, since in all our expressions we require the back-action of the electromagnetic field emitted by the particle onto itself (via scattering by the material bodies), source and detection lateral positions coincide and we retain in the Green tensor only one spatial argument. In addition, given the system’s translational invariance along the particle trajectory, we carry out a Fourier transform, so that the Green tensor acquires a mixed Fourier- and real-space dependence with regards to the positional arguments and, of course, features a dependence on frequency, i.e. $\underline{G}(q,\mathbf{R}_{a},\omega)$. This expression fulfills the identities
@@
which we may, respecitively, view as consequences of the system’s reciprocity in the real space and stationarity in time.
@@
This would lead to an expression of the same form as provided in Eq. (21). However, due to the (system-specific) condition in Eq. (25a), we can go one step further and find that the coefficients associated with $\boldsymbol{\underline{\Lambda}}_{\perp}$ and $\underline{L}_{n}$ actually vanish (see App. B) so that we can write
@@
$\mathrm{Tr}\left[\underline{S}^{\sf T}\underline{G}_{\Im}\right]=\mathrm{Tr}\left[\underline{\Delta}\underline{D}\right]+2\chi_{n}c_{n}-2\boldsymbol{\Omega}_{\perp}\cdot\mathbf{s}_{\perp}.$ (28a)
@@
$\stackrel{{\scriptstyle\text{average}}}{{=}}\frac{d^{2}}{3}\mathrm{Tr}\left[\underline{G}_{\Im}\right]\;\Sigma=\frac{d^{2}}{3}\mathrm{Tr}\left[\underline{D}\right]\;\Sigma~.$ (28b)
@@
The last line in the above equations corresponds to the result of an average over all possible directions of the vector $\mathbf{d}$ for which the power-spectrum tensor becomes proportional to $\underline{\mathds{1}}d^{2}/3$ ($d^{2}=\mathbf{d}\cdot\mathbf{d}$)  [21, 92, 77]. When comparing the two results, we notice, as expected, that the main difference lies in the contribution arising from $\boldsymbol{\Omega}_{\perp}$, which is absent in the second model. Moreover, when $\boldsymbol{\Omega}_{\perp}$ and $\mathbf{s}_{\perp}$ have the same sign, the corresponding contribution tends to reduce the frictional force. This confirms and generalizes earlier results in Refs. [22, 73], showing that they are not tied to the specific model for the dipole’s dynamics used there (see Sec. 3.2). The remaining contributions in Eqs. (28) are similar in both models although not identical. Specifically, the terms proportional to $c_{n}$ directly show how the system’s symmetry (or better the lack thereof) affects the frictional force. More precisely, as soon as the system is mirror-symmetric with respect to a plane containing the particle’s trajectory – for example when the particle moves above a planar structure – the terms proportional to $c_{n}$ in Eqs. (28) vanish. Also, averaging over the dipole direction in the second model, washes out this contribution which does not appear in the last line of Eq. (28).
@@
Another interesting consequence is that the results in Eqs. (28) are both equivalent to the identity
$\mathrm{Tr}\left[\underline{S}^{\sf T}(-\omega^{+}_{q},-v)\underline{G}_{\Im}(q,\mathbf{R}_{a},\omega)\right]=\mathrm{Tr}\left[\underline{S}(-\omega^{+}_{q},v)\underline{G}_{\Im}(q,\mathbf{R}_{a},\omega)\right]~.$ (29)
@@
When used in Eq. (24) in conjunction with Eq. (25a) and after changing $q\to-q$, the previous expression allows us to show that, in general, the frictional force must always reverse its sign if $v\to-v$, i.e. $F_{\rm fr}(-v)=-F_{\rm fr}(v)$. In other words, independently of the model for the dipole’s dynamics, $F_{\rm fr}(v)$ is an odd function of the velocity. This is what is physically expected for a frictional interaction and has been confirmed by calculations with specific models [93, 94, 95, 67, 71, 72]. We want to underline, however, that the mathematical properties justifying this behavior are less trivial than one could intuitively presume. Indeed, reversing the logic and imposing that $F_{\rm fr}(v)$ is odd with respect to the velocity one would obtain Eq. (29), which can erroneously suggest that $\underline{S}^{\sf T}(\omega,-v)=\underline{S}(\omega,v)$, in contradiction with Eq. (21).
3.2 An illustrative example
@@
Thus far, we have refrained from assuming particular microscopic dynamics for the electric dipole operator. To illustrate the validity and physical implications of our results, we now introduce a tractable model that admits an exact solution for the equations of motion. Specifically, we assume that the particle moving in vacuum parallel to a translational invariant arrangement of bodies is an atom. Given that the interaction between the atomic internal degrees of freedom and the electromagnetic fluctuations is typically weak, it is reasonable to describe the dipole operator in term of linear system associated with a single frequency $\omega_{a}$. Physically, $\omega_{a}$ would correspond to the transition frequency between the highest occupied orbital to the lowest unoccupied one. Within this linear approximation, we are going to represent $\hat{\mathbf{d}}$ in terms of a 3D isotropic quantum-mechanical oscillator  [96, 67] which fulfills the following equation of motion
@@
[…]
@@
With the above results, we can explicitly check the general properties of the power-spectrum tensor which we have discussed in Sec. 2 and at the beginning of Sec. 3. To start, we notice that $\underline{S}$ is Hermitian and in general not real-valued, as expected. It does not fulfill the FDT but, instead, fulfills Eq. (14). To see this more clearly, we consider the following identity which we can derive from Eq. (34)
@@
$\times\underline{\alpha}(\omega,v)\underline{G}_{\Im}(q,\mathbf{R}_{a},\omega_{q}^{+})\underline{\alpha}^{{\dagger}}(\omega,v)~.$ (38)
Upon utilizing the properties of the polarizability and of the Green tensors mentioned above, we can show that, consistent with the discussion in Sec. 2, $\underline{J}$ is Hermitian and fulfills the crossing relation. Moreover, in earlier work with our co-workers, we have demonstrated within the above framework that the local thermal equilibrium approximation displays significant drawbacks regarding the quantitative evaluation of quantum friction. Specifically, neglecting $\underline{J}$, i.e. enforcing the local thermal equilibrium approximation, underestimates the strength of the interaction and adversely affects the thermodynamical consistency of the physical system  [21, 22, 23].
Consider now the power-spectrum tensor. Noting that $(1-e^{x})^{-1}=1-(1-e^{-x})^{-1}$ and utilizing the expressions in Eqs. (25) and (37) as well as the properties of the dressed polarizability, we find
@@
With the above, we can already check that the expression in Eq. (40) has a form that is consistent with Eq. (27) where $c_{n}=0$ and $\mathbf{s}_{\perp}=s_{y}\mathbf{y}$. As discussed earlier, we may view this as being the consequence of the mirror-symmetry of the total system with respect to the $y=0$ plane. Finally, inserting in the previous equations the expression for the Green tensor given above and using Eq. (35), one can show that $\underline{S}(\omega,v)$ is given by the sum of a diagonal and an anti-diagonal matrix. This is in agreement with our general considerations about the structure of the tensor based on the symmetry of the system. Moreover, we can verify that $\underline{S}(\omega,-v)\neq\underline{S}^{\sf T}(\omega,v)$, as we should expect in terms of our general result in Eq. (21). Notice that this result is obtained despite the fact that $\underline{\alpha}(\omega,-v)=\underline{\alpha}^{\sf T}(\omega,v)$ and that from Eq. (36) $\underline{\mathcal{D}}(\omega,-v)=\underline{\mathcal{D}}^{\sf T}(\omega,v)$. We can further verify that
@@
where, in agreement with the relation in Eq. (21), $f(\omega,v)$ is a real function odd in $v$. The previous expression also implicitly confirms that the diagonal elements of the power-spectrum tensor are even in $v$, while the off-diagonal terms arising from $\boldsymbol{\Omega}_{\perp}=\Omega_{y}\mathbf{y}$ are odd in the same variable.
4 Conclusions
@@
In summary, we have discussed a number of relations involving the power-spectral density of a, in general, quantum system in an arbitrary steady state. This includes a large class of far-from-equilibrium scenarios as well as the equilibrium situation as a limiting case. We have adopted a minimalistic viewpoint and, besides stationarity, our analyses do not rely on any specific information about the microscopic dynamics of the system. Therefore, our results have broad validity, including systems with non-Markorvian as well as nonlinear dynamics. This is particularly relevant for quantum systems in which the Markov approximation may lead to inconsistent results [4, 51].
@@
More precisely, we have focused on the steady state of a multidimensional (multivariate) stochastic process, where the dynamics is described in terms of a three-dimensional vector operator. Since the steady state may correspond to nonequilibrium dynamics for which the principle of detailed balance can be violated, close-to-equilibrium relations such as the fluctuation-dissipation theorem are not necessarily fulfilled. Nonetheless, we have been able to derive and classify a collection of properties that characterize the power spectrum and its connection to other physical quantities such as the system’s susceptibility. When appropriate, we have linked them to the system’s dynamics, its quantum behavior, the flow of energy within it and the role of dissipation, discussing their close-to-equilibrium counterparts. This provides us with a solid basis to investigate and assess the validity of approaches such as the local thermal equilibrium approximation.
@@
Once additional information about the system, such as symmetries and details about its structure, is provided, we may obtain further properties and relations among the different quantities that characterize the system. As an illustration, we have analyzed the nonconservative viscous dynamics of a particle that moves in the quantum electromagnetic vacuum parallel to a translationally invariant arrangement of material bodies. When the drag felt by the particle during its motion is balanced by an external driving force, the system reaches a nonequilibrium steady state that is characterized by a speed $v$. Without specifying the internal microscopic dynamics of the moving object, we have obtained a collection of relations that characterize the power-spectrum tensor, its dependence on the velocity $v$ and its role in determining the frictional interaction at zero temperature (quantum friction). In addition, for this system, we have tested the general results described in the main part of the manuscript, in the context of a specific model which provides the microscopic internal dynamics of a moving atom. In particular, we have modeled the atom’s dipole vector operator in terms of a three-dimensional harmonic oscillator and have derived and examined the corresponding power-spectrum tensor as well as the related matrices.
@@
Our results can be useful in a large variety of systems and specifically for numerous fluctuation-induced phenomena, such as the (quantum) Brownian motion [100, 101, 102], radiative heat-transfer [103, 83, 104, 105] the Casimir effect [106, 107, 108] and the Casimir-Polder interaction [109, 110, 111], including their nonequilbrium variants such as quantum friction [15, 43, 44, 45]. Since often the information about the statistical properties of the system is or can be encoded in the power spectral density, knowing its general properties can be very helpful for understanding its behavior and to test thermodynamic consistency of its descriptions [23].